### v1

# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

Abstract
Nonlocal Coulomb interactions play a crucial role in stabilizing distinct electronic phases in kagome materials. In this work, we systematically investigate the effects of on-site ($U$) and inter-site ($V$) Coulomb interactions on the electronic structure and stability of charge-density-wave (CDW) phases in the kagome metal KV3Sb5 using density functional theory (DFT+$U$+$V$) calculations. We demonstrate that $V$ promotes the formation and stability of CDW phases, whereas $U$ suppresses these phases, highlighting a fundamental competition between local and nonlocal electronic correlations. By directly comparing our theoretical results with angle-resolved photoemission spectroscopy (ARPES) data, we identify realistic values of $U$ and $V$ that accurately describe the electronic band structure of KV3Sb5. Our findings establish a detailed $U$$V$ phase diagram for KV3Sb5, offering valuable insights into the correlated electronic states in kagome metals and serving as a foundation for future explorations of correlation-driven phenomena in related materials.
preprint: APS/123-QED
I Introduction
Kagome lattices are formed by a periodic arrangement of corner-sharing triangles, resulting in a distinctive crystal structure with inherent geometric frustration Freedman et al. (2010). This geometric frustration arises from the nonequivalence of the three vertices in each triangular unit, which prevents the system from achieving long-range magnetic ordering and instead stabilizes disordered spin states. Consequently, kagome lattices with intrinsic magnetic frustration are promising candidates for hosting quantum spin liquid states Balents (2010); Norman (2016); Georgopoulou et al. (2023); Han et al. (2012).
In addition, kagome lattices exhibit topologically nontrivial electronic states linked to local electronic symmetry breaking Ye et al. (2018); Yin et al. (2018); Liu et al. (2020); Kang et al. (2020); Morali et al. (2019). Their complex band structures—featuring flat bands, Dirac cones, and van Hove singularities—heighten their significance in condensed matter physics Yin et al. (2022). The interplay between strong Coulomb interactions and nontrivial band topology, driven by the diverging density of states at van Hove singularities and flat bands, gives rise to diverse exotic phenomena. These include topological insulator Guo and Franz (2009), superconductivity Ko et al. (2009), charge density wave (CDW) Wang et al. (2013), charge fractionalization Rüegg and Fiete (2011); Mizoguchi et al. (2017), and Dirac/Weyl semimetals Morali et al. (2019); Liu et al. (2019).
[…]
Regarding their crystal structures, initial results from X-ray diffraction Ortiz et al. (2020, 2021b) and scanning tunneling microscopy (STM) Jiang et al. (2021); Zhao et al. (2021); Liang et al. (2021a) indicate the presence of a $2\times 2$ CDW phase. A recent density functional theory (DFT) study, in comparison with STM data, confirmed the inverse Star of David (ISD) distortion as the ground state for AV3Sb5 compounds Tan et al. (2021). Subsequent studies have provided further evidence supporting the ISD deformation in these systems Kato et al. (2022); Kang et al. (2022).
Important features of the electronic structures in $A$V3Sb5 compounds are the van Hove singularities near the Fermi energy, which have V-$3d$ orbital character. The precise position of these van Hove singularities depends on details of the Coulomb interactions, inducing electronic instabilities that lead to the emergence of various CDW phases Jiang et al. (2021); Denner et al. (2021); Jeong et al. (2022). In particular, the chiral CDW phase observed in $A$V3Sb5 has been plausibly attributed to nonlocal Coulomb interactions Denner et al. (2021). Other contributing factors, such as lattice deformations, electron-phonon coupling, and Fermi surface nesting, can also interplay with electronic correlations and play significant roles in mediating the CDW phase transition Tan et al. (2021); Wenzel et al. (2022); Xie et al. (2022); Ratcliff et al. (2021); Luo et al. (2022); Wu et al. (2022); Liu et al. (2022).
The impact of Coulomb interactions has been extensively investigated using Kagome Hubbard model calculations, primarily within single-band models. Kiesel et al. Kiesel and Thomale (2012); Kiesel et al. (2013) explored the system using a single-band extended Hubbard model that incorporates both on-site ($U$) and inter-site ($V$) Coulomb interactions. They studied various phases arising from Fermi surface instabilities within the $U$-$V$ phase diagram Kiesel and Thomale (2012); Kiesel et al. (2013). Notably, near the van Hove filling, $V$ can promote distinct CDW phases, which explain several key experimental findings Ferrari et al. (2022). The presence of sizable long-range interactions has been also reported, which exhibit interesting universality across kagome metals Di Sante et al. (2023). More recent studies have further emphasized the critical role of inter-site Coulomb interactions in understanding the electronic structure of kagome materials, investigating their impact on a broad range of electronic properties Wu et al. (2021); Zheng et al. (2022); Kiran et al. (2024); Jiang et al. (2025); Dong et al. (2023); Bigi et al. (2024); Wenger et al. (2024).
In this study, we perform DFT calculations, incorporating both $U$ and $V$ Coulomb interactions to explore their specific roles in the electronic structure of KV3Sb5. We examine the energy profiles associated with in-plane distortions, leading to the Star of David (SD) and ISD phases, along with the variations in $U$ and $V$. We also explore the effects of $U$ and $V$ on different stacking arrangements along the out-of-plane direction of various CDW phases and present the resulting $U$-$V$ phase diagram for KV3Sb5. A comparison with the experimental ARPES data and our calculations enables us to locate the position of the system in the phase space.
Figure 1: (a) Unit cell and (b) $2\times 2\times 1$ supercell crystal structure of KV3Sb5, where the vanadium atoms exhibit kagome net and the V-Sb layers are intercalated by a layer of K atoms. (c) Star of David (SD) and (d) inverse Star of David (ISD) $2\times 2$ charge order distortion in KV3Sb5. The K, V, Sb atoms are represented by brown, dark blue, and light blue balls, respectively. V1 and V2 represent the inequivalent vanadium sites of the kagome lattice.
Figure 2: (a) Variation of energy profiles as a function of the total distortions per formula unit (f.u.) for KV3Sb5: (a-d) for fixed $U=$ 0, 1, 2, and 3 eV with varying $V$, and (e-h) for fixed $V=$ 0, 1, 2, and 3 eV with varying $U$. The total distortion refers to the sum of the displacements of all vanadium atoms within the $2\times 2\times 1$ supercell. $\Delta E$ represents the energy difference relative to the pristine phase. The positive (negative) distortion corresponds to the magnitude of the breathing (shrinking) phonon mode of the $2\times 2$ inverse Star of David (Star of David) CDW phase.
II Computational details
Density functional theory (DFT) calculations were performed using the plane-wave pseudopotential method, as implemented in the Quantum Espresso package Giannozzi et al. (2009, 2017a, 2017b). The exchange-correlation functional was treated using the generalized gradient approximation (GGA) parameterized by Perdew-Burke-Ernzerhof (PBE) Perdew et al. (1996). Pseudopotentials were obtained from the Standard Solid-State Pseudopotentials (SSSP) library, version 1.2.1 SSS . We considered a $2\times 2\times 1$ supercell of KV3Sb5, containing 36 atoms, to incorporate two relevant CDW phases--SD and ISD--on equal footing (see Fig. 1), and performed all calculations using a $5\times 5\times 5$ $k$-mesh. The plane-wave energy cut-off was set to 37 Ry. For density of states calculations, a denser $8\times 8\times 8$ $k$-mesh was used. Band unfolding calculations were carried out using the unfold.x code Pacilè et al. (2021), based on the method of Popescu and Zunger Popescu and Zunger (2012). Spin-orbit coupling and van der Waals corrections were not included, as they do not affect the main conclusions of this study.
The optimized structure obtained from the DFT calculations yields lattice parameters of $a=5.48$ and $c=9.35$. These optimized lattice parameters are in good agreement with experimental data Ortiz et al. (2019, 2021b) and previous DFT studies Tan et al. (2021); Ortiz et al. (2021b). Neglecting van der Waals corrections results in a slightly larger $c$ lattice parameter compared to experimental reports ($a_{\text{expt}}=5.48$, $c_{\text{expt}}=8.958$) Ortiz et al. (2021b, 2019). Details of the structural parameters are provided in Table S1 of the Supplementary Material Sup .
The Coulomb interaction parameters ($U$ and $V$) for the V-$3d$ manifold were calculated using the constrained random-phase approximation (cRPA) with the projector method Kaltak (2015) implemented in VASP Kresse and Furthmüller (1996). A $8\times 8\times 5$ $k$-mesh and 288 bands were employed for the cRPA calculations. To construct the Wannier functions, all V-$3d$ and Sb-$5p$ orbitals were considered, with a disentanglement energy window of [-5.0, 4.8] eV (see Fig. S1). The Wannier functions were obtained using WANNIER90 Mostofi et al. (2008) and the VASP2WANNIER Franchini et al. (2012) interface, with default parameters Reddy et al. (2024) for disentanglement and Wannierization. A large plane-wave cutoff energy of 500 eV was used for self-consistent calculations, while 400 eV was used for dielectric function calculations. The resulting band structures and interaction parameters are provided in the Supplementary Material Sup . Using the $U$ and $V$ values obtained from cRPA calculations, we employed the DFT+$U$+$V$ method, as implemented in Quantum Espresso Campo and Cococcioni (2010), to investigate the electronic structure of KV3Sb5. In addition, we systematically varied the values of $U$ and $V$ to further investigate their impact on the ground state and electronic structure of KV3Sb5.
III Results and Discussion
III.1 Pristine and CDW structures
Figures 1(a) and (b) show the unit cell and $2\times 2$ supercell of KV3Sb5, respectively, with space group P6/$mmm$ (No. 191). The supercell contains two inequivalent vanadium sites, denoted as V1 and V2, whereas the unit cell contains only one. The vanadium sublattice forms a perfect kagome lattice structure, with the V-Sb layers inserted between the K atoms. Two different $2\times 2$ CDW structures, SD (Fig. 1(c)) and ISD (Fig. 1(d)), were constructed by incorporating opposite lattice distortions on the V atoms relative to the pristine supercell.
III.2 Stability of CDW phases depending on $U$ and $V$
First, we discuss the stability of the CDW phases upon introducing $U$ and $V$ in the kagome metal KV3Sb5. Figure 2 shows the energy profiles as a function of vanadium atomic distortions for different values of $U$ and $V$. Positive (negative) distortion values correspond to the ISD (SD) phase. Our results clearly demonstrate that, in the absence of $U$ and $V$, the ISD deformation is energetically more favorable than the SD deformation. This is consistent with experimental STM images and previous DFT results, where the ISD phase is found to be more stable than the SD phase Tan et al. (2021); Jiang et al. (2021). For all fixed values of $U$, increasing $V$ not only enhances the stability of both the ISD and SD phases but also increases atomic distortions (see Figs. 2(a-d)). The ISD phase remains more stable than the SD phase upon variation of $V$. For instance, at $U$ = 0, the atomic distortion increases by a factor of 2.9 (ISD phase) and 3.2 (SD phase) for $V=3$ eV compared to the case of $V=0$. Furthermore, the energy difference between the ISD and SD phases reaches 204 meV, a significant increase from 8 meV at $U$ = 0 and $V=0$.
Interestingly, for a fixed $V$, increasing $U$ progressively reduces the stability of both the ISD and SD phases. Eventually, at $V=0$ and $U=3$ eV, the pristine phase becomes energetically more favorable than the CDW phases (see Fig. 2(e)). This demonstrates that strong local Coulomb interaction alone cannot stabilize either of the CDW phases. For finite values of $V$, higher $U$ values ($\geq 3$ eV) are required to stabilize the pristine phase relative to the CDW phases. These results indicate that $V$ and $U$ exert opposite effects: while $V$ stabilizes CDW phases, $U$ suppresses them. This contrast provides compelling evidence that incorporating both $U$ and $V$ leads to fundamentally different physical consequences in the kagome metal KV3Sb5.
We compare these findings for KV3Sb5 with those for another kagome metal, CsV3Sb5. Employing the similar approaches, we obtained the corresponding energy profile for CsV3Sb5 as shown in Fig. S2 of the Supplementary Material Sup . For a fixed $U$, CsV3Sb5 follows a similar trend to KV3Sb5, except when $V=0$. When $V=0$, as $U$ increases (for $U<3$ eV), the ISD phase gradually destabilizes prior to the SD phase. At $U=3$ eV, the SD phase also becomes destabilized, approaching the pristine phase. These results shows common tendency toward ISD phase for both systems. Note that the effective values of $U$ and $V$ in CsV3Sb5 are smaller than in KV3Sb5 suggesting the stronger CDW instability of the system.
Figure 3: Phase diagram of KV3Sb5 of depending on $U$ and $V$. The symbols indicate the values of $U$ and $V$ for which the calculations have been performed. The filled blue, red, and green circle symbols represent the ISD+ISD ($\pi$-shift), ISD+ISD (zero-shift), and pristine phase, respectively. The yellow star symbol represents the $U$ and $V$ values obtained from cRPA.
III.3 Phase diagram of KV3Sb5
Previous studies on the $A$V3Sb5 system have demonstrated the possibility of various CDW phases, which not only differentiate in-plane SD and ISD configurations but also include phase shifts between adjacent kagome layers Liang et al. (2021a); Li et al. (2021); Ortiz et al. (2021a); Tan et al. (2021) (see Fig. S3). Motivated by these findings, we considered six distinct CDW phases by employing the SD and ISD deformations with either zero or $\pi$ phase shifts between adjacent kagome layers in a $2\times 2\times 2$ supercell. These six CDW phases include (a) ISD+ISD ($\pi$-shift), (b) ISD+ISD (zero-shift), (c) SD+SD ($\pi$-shift), (d) SD+SD (zero-shift), (e) SD+ISD (zero-shift), and (f) SD+ISD ($\pi$-shift). Note that the ISD+ISD with zero phase shift is equivalent to the $2\times 2\times 1$ ISD structure discussed in Sec. III.2.
Figure 3 shows the overall CDW phase diagram of KV3Sb5 as a function of $U$ and $V$. We first note that within the studied $U$-$V$ parameter range, the ISD deformation is consistently favored, as shown in Fig. 2. This preference arises because the energy scale of the inter-layer coupling is significantly smaller than that of the intra-layer CDW distortion. Except in the limit of large $U$, where the pristine phase is stabilized for small $V$, ISD-related phases remain more stable than SD phases throughout the phase diagram.
Among the various phases, two dominant ISD-type CDW phases, ISD+ISD ($\pi$-shift) and ISD+ISD (zero-shift), are noted in the phase diagram. We found that in the small $U$ and $V$ regime, ISD+ISD ($\pi$-shift) is stabilized, as shown in Fig. 3. We find that inter-site Coulomb interactions promote coherent ISD deformations across different kagome layers, favoring a zero-shift phase alignment. For sufficiently large $V$, ISD+ISD (zero-shift) is always stabilized regardless of $U$. However, in the small $V$ regime, competition arises between ISD+ISD ($\pi$-shift) and either ISD+ISD (zero-shift) or the pristine phase, depending on whether $U$ is small or large. This finding is intriguing because our Coulomb interaction profiles are restricted to the in-plane interactions, yet they selectively stabilize specific inter-layer phase shifts. It also highlights that competing phases can be stabilized by subtle details of the electronic structure.
Previous studies have identified the ISD+ISD ($\pi$-shift) phase as the ground state of the $A$V3Sb5 system Liang et al. (2021b); Jin et al. (2024); Li et al. (2022). Hence, we can determine the realistic regime of the on-site $U$ and inter-site $V$ Hubbard parameters for the KV3Sb5 system in the $U$-$V$ phase diagram: small to moderate $U$ values and small $V$ values (see the blue-filled areas in Fig. 3). To further quantify the Coumonb interaction parameters, we performed cRPA calculations, obtaining $U$ = 1.35 eV and $V$ = 0.17 eV. The corresponding point is marked with a star symbol in Fig. 3, demonstrating that KV3Sb5 lies well within the ISD+ISD ($\pi$-shift) regime, consistent with experimental findings. The validity of these $U$ and $V$ values obtained from cRPA will be further examined in the next section through direct comparison between the band dispersion from DFT+$U$+$V$ calculations and experimental ARPES data.
III.4 Electronic structures: band structure
Figure 4: Calculated unfolded band dispersion for $2\times 2$ ISD phase along K-M-K path for different values of $U$ and $V$ (in eV): (a) cRPA values ($U=1.35,V=0.17$), (b) $U=0$, $V=0$, (c) $U=1$, $V=0$, (d) $U=1.5$, $V=0$, (f) $U=1.5$, $V=0.5$, (g) $U=1.5$, $V=1$, and (h) $U=1.5$, $V=1.5$. (e) ARPES intensity measured at $T$ = 20 K along the K-M-K cut at $h\nu=114$ eV, as adapted from Kato et al. Kato et al. (2022). The black solid lines in (a) are a guide for the eye to understand both the SP1 and SP2 bands. A Fermi level offset of 200 meV is introduced in the DFT+$U$+$V$ bands to enable direct comparison with the experimental spectra.
We now investigate the electronic structure of KV3S5, focusing on its dependence on the Hubbard parameters $U$ and $V$. Given the weak interlayer interactions, we consider the $2\times 2$ ISD phase, as shown in Fig. 1. To isolate the effects of $U$ and $V$ from structural contributions, we use the optimized ISD structure obtained at $U=0$ and $V=0$ as the baseline for calculations with various $U$ and $V$ values. The phase shifts between adjacent kagome layers have minimal impact on the band dispersion. Thus, our approach is sufficient for investigating the electronic structure of KV3S5. Structural contributions, such as variations in vanadium atom distortion for different values of $U$ and $V$, will be discussed at the end of this subsection.
Figure 4 presents the band dispersion of the $2\times 2$ ISD phase for various values of $U$ and $V$, unfolded onto the $1\times 1$ Brillouin zone. The $k$-path K-M-K is selected to facilitate direct comparison with existing angle-resolved photoemission spectroscopy (ARPES) measurements, as shown in Fig. 4(e).
The band dispersion of the ISD phase successfully captures two dominant features near the Fermi level: (i) an M-shaped band and (ii) a $\Lambda$-shaped band, as highlighted by the guidelines in Fig. 4(a). These features are consistent with the experimental ARPES measurement by Kato et al. Kato et al. (2022), where the former is denoted as the SP1 band and the latter as SP2. Near the M-point, the $\Lambda$-shaped (SP2) band penetrates the M-shaped (SP1) band and reaches a higher binding energy in the ISD phase. In contrast, no such penetration occurs in the SD phase, where the $\Lambda$-shaped (SP2) band remains at a higher energy position.
We found that our on-site and inter-site Hubbard parameters, $U$ = 1.35 eV and $V$ = 0.17 eV, obtained from the cRPA calculations, successfully reproduce the key features of the experimental ARPES described above (Fig. 4 (a)). Both the relative energy positions of the SP1 and SP2 bands and their morphology closely match the experimental spectra in Fig. 4(e). This agreement contrasts with the results obtained using other $U$ and $V$ values, as shown in the Fig. 4. While DFT calculations without $U$ and $V$ (Fig. 4(b)) also show reasonable agreement with ARPES data, the cRPA-derived finite $U$ and $V$ values play a crucial role in shaping the band dispersion of the KV3Sb5 system, as demonstrated in Fig. S5 of the Supplementary Material Sup .
Our systematic investigation reveals the distinct roles of $U$ and $V$ in shaping the low-energy band structures. Variations in $U$ primarily affect the M-shaped (SP1) band, which predominantly has $3d_{x^{2}-y^{2}}$, $3d_{xy}$, and $3d_{z^{2}}$ orbital characters,
while the $\Lambda$-shaped (SP2) band, mainly associated with $3d_{xz/yz}$ orbital characters, remains nearly rigid (see Fig. S7 in the Supplementary Material Sup for detailed orbital characters). Note that the orbital characters of both SP1 and SP2 bands are consistent with the earlier findings by Kato et al. Kato et al. (2022). Increasing $U$ progressively shifts the SP1 band downward, lowering its energy at the $K$-point, whereas $U$ has minimal impact on the position of the SP2 band. In contrast, the effect of $V$ is far more pronounced. Fixing $U$ at 1.5 eV and varying $V$ from 0.5 to 1.5 eV, we find that the low-energy band structure is highly sensitive to $V$. As $V$ increases, the M-shaped nature of the SP1 band is completely disrupted, and the relative position of the SP1 and SP2 bands changes significantly.
The band dispersion of the ISD phase across the entire Brillouin zone (calculated using the cRPA-derived values: $U$ = 1.35 eV and $V$ = 0.17 eV), is provided in Fig. S6 Sup , along with that of the pristine phase for comparison.
The positions of the van Hove singularities are essential for identifying electronic instabilities Li et al. (2021); Ferrari et al. (2022); Xie et al. (2022). In the pristine phase, one of the prominent van Hove singularities appears near the M-point. This feature is found to be split in the ISD phase, as demonstrated in Fig. S6 Sup . The split bands are associated with the M-shaped band shown in Fig. 4(a). The dominant orbital contributions to the M-shaped band arise from the $3d_{x^{2}-y^{2}}$, $3d_{xy}$, and $3d_{z^{2}}$ orbitals, which give rise to pronounced peaks near -0.2 eV in the density of states (see Fig. 5). Note that the van Hove singularity peaks are substantially suppressed in the ISD phase relative to the pristine phase, due to the band splitting near the M-point.
Now We account for structural contributions corresponding to different $U$ and $V$ values, as discussed in Sec. III.2, which were neglected in the previous calculations. Figure S5 (in the Supplementary Material Sup ) shows the band-unfolding results for the optimized $2\times 2$ ISD structures with corresponding $U$ and $V$ values. For higher $V$ values, structural optimization induces large atomic displacements of vanadium atoms, resulting in significant energy band splitting. Notably, in the absence of $V$, the M-shaped nature of the SP1 band is almost entirely suppressed as $U$ increases from 0 to 2 eV. This suggests that the unfolded band dispersion closely resembles that of the pristine structure (shown in Fig. S4 Sup ), reinforcing our earlier analysis that CDW deformations diminish and the system transitions towards a stable pristine structure as $U$ increases. It is worth noting that the primary effect of the structural contributions is the modification of the band dispersion arising from the larger atomic distortions at finite $U$ and $V$ values.
III.5 Electronic structures: density of states
Figure 5: (a) and (b) Orbital-resolved density of states for two different vanadium sites near the Fermi level in the $2\times 2$ ISD phase, calculated using $U$ and $V$ obtained from cRPA. The Fermi level is set to 0 eV. (c) Calculated V-$3d$ occupation numbers for the $2\times 2$ ISD phase obtained for different $U$ and $V$ values. The filled circles and cross symbols represent the two inequivalent vanadium sites V1 and V2, respectively. (d) Top view of the kagome lattice plane with its respective local (x, y) axes. The local x-axis points toward the Sb atom in the kagome lattice plane, while the local y-axis points to the center of the vanadium triangle.
To gain deeper insight into the electronic structure of KV3Sb5, we calculated the density of states (DOS) based on the local axis, as shown in Fig. 5(d). In the $2\times 2$ ISD phase, vanadium atoms occupy two inequivalent sites in the kagome lattice, denoted as V1 and V2 (Fig. 1). The orbital-resolved DOS for V1 and V2 in the $2\times 2$ ISD phase is shown in Fig. 5. To examine the effects of $U$ and $V$ solely, we used the optimized ISD structure obtained at $U=0$ and $V=0$. The DOS for V1 and V2 exhibit similar behavior, with V1 contributing slightly more than V2 near the Fermi level, which remains unchanged for all values of $U$ and $V$. Furthermore, in the absence of $V$, increasing $U$ enhances the DOS at the Fermi level, whereas incorporating $V$ (without $U$) leads to the opposite trend. Specifically, $V$ suppresses the DOS at the Fermi level, except for the $d_{xz/yz}$ orbitals, which contribute to the $\Lambda$-shaped (SP2) band, as shown in the projected band dispersion in Fig. S7 Sup . This interplay between $U$ and $V$ in shaping the PDOS at the Fermi level is illustrated in Fig. S8 of the Supplementary Material Sup . It is well established that CDW phases disrupt van Hove singularities, leading to a reduction in the DOS at the Fermi level Tan et al. (2021). These finding align with our earlier results in Sec. III.2, where we observed that increasing $U$ destabilizes the CDW phase, making the pristine structure more favorable, while higher values of $V$ help sustain the CDW phase.
We further calculated the $3d$ occupation numbers for V1 and V2 across different values of $U$ and $V$. These were determined by integrating the DOS up to the Fermi energy in the valence band. Figure 5(c) shows the occupation numbers for V1 and V2 as a function of $U$ for different $V$ values in the $2\times 2$ ISD phase. For $U=0$ and $V=0$, the occupation number of V1 (V2) is nearly 4.102 (4.093), decreasing to 4.055 (4.056) at $U=3$ eV and $V=0$. Our results confirm that for any given $V$, the occupation number decreases with increasing $U$, whereas for a fixed $U$, increasing $V$ leads to the opposite trend. The orbital-decomposed occupation numbers, shown in Fig. S9 Sup , further illustrate this contrasting behavior of $U$ and $V$.
IV Conclusions
In summary, we systematically investigated the roles of on-site ($U$) and inter-site ($V$) Coulomb interactions in shaping the charge-density-wave (CDW) phases of the kagome metal KV3Sb5. From our DFT+$U$+$V$ calculations, we constructed the $U$-$V$ phase diagram, revealing the contrasting roles of $U$ and $V$ in forming CDW phases. We found that $V$ plays a decisive role in stabilizing the inverse Star of David (ISD) CDW phase, consistent with experimental observations. The interplay between $U$ and $V$ influences not only the relative stability and extent of lattice distortions but also modulates the electronic instabilities associated with van Hove singularities near the Fermi level. By examining unfolded band structures, we clearly demonstrated how variations in $U$ and $V$ modify the band shapes and dispersions. Our cRPA estimates provided realistic values of $U$ and $V$, enabling us to identify the position of KV3Sb5 in the $U$-$V$ phase space. Through comparison with experimental ARPES data, we demonstrated that our chosen interaction parameters accurately reproduce the band structure of the system.
Our findings underscore the importance of both local and nonlocal electronic correlations in kagome systems and provide a unified framework for understanding the emergence of exotic phases—including unconventional superconductivity and topological phenomena—in kagome metals. This work offers valuable insights for future studies of correlated states in kagome-lattice materials.
Acknowledgements
BK acknowledges support from NRF (NRF2021R1A4A1031920, RS-2021-NR061400, and RS2022-NR068223) and KISTI Supercomputing Center (Project No. KSC-2023-CRE-0413). CJK was supported by the National Research Foundation of Korea (NRF) grant funded by the Korean Government (MSIT) (Grant No. 2022R1C1C1008200).